# CogFlow: Bridging Perception and Reasoning through Knowledge Internalization for Visual Mathematical Problem Solving

### v1

@@ Appendix B Additional Details of MathCog Dataset
Existing multimodal reasoning corpora provide abundant natural language annotations but rarely disentangle perception from reasoning (zhang2024mavismathematicalvisualinstruction; gao2023gllavasolvinggeometricproblem). This gap makes it difficult to supervise models in a way that enforces accurate perception and strengthens their ability to internalize visual content. To address this, we construct the MathCog dataset, which explicitly separates the watching (perception) and thinking (reasoning) stages. Furthermore, to support the different training phases of CogFlow, we curate three tailored subsets: MathCog-SFT, MathCog-IntlzR, and MathCog-RL.
@@ B.2.1 Data Collection
To construct MathCog, we first collect a large pool of geometry-related visual math problems from existing corpora, including MAVIS (zhang2024mavismathematicalvisualinstruction), Geo170K (gao2023gllavasolvinggeometricproblem), and LLaVA-CoT (xu2025llavacotletvisionlanguage). We then perform a careful filtering process to obtain a high-quality subset of 111,752 problems. Specifically, we remove (1) problems dominated by weak geometric relevance charts (e.g., histograms, pie charts) or artistic illustrations that deviate from geometric reasoning, and (2) diagrams with either excessively low resolution (<< 64) or overly high resolution (>> 2048), which would otherwise compromise model training consistency.
@@ B.2.2 MathCog-SFT and MathCog-RL Curation
To unify the representation, the extracted elements are normalized and remapped to a fixed coordinate system, following the normalization scheme in prior work (wei2025slowperceptionletsperceive):
@@ B.2.3 MathCog-IntlzR Curation
We first sample 10,000 examples as positive samples from MathCog-SFT for MathCog-IntlzR. As mentioned earlier, we categorize the errors into five types: (1) omitting or misbinding primitives, (2) introducing nonexistent facts, (3) contradicting geometric constraints, (4) invoking external theorems inappropriately, and (5) referring inconsistently to established elements. The negative data for MathCog-IntlzR is constructed based on these five error types. The curation of MathCog-IntlzR is illustrated at the bottom of the Figure 10. We first use an LLM, such as GPT-5 (gpt5), to generate negative trajectories conditioned on each positive trajectory by explicitly instructing it to modify the corresponding “Thinking” part, as specified in the prompt shown in Table 7. The generated candidates are then manually checked. For example, for the Type 1 error: Omit or misbind primitives in Figure 11, we input the designed prompt into the LLM and obtain the variant “In any parallelogram, if the perimeter of triangle CDM is 5, …”, which deliberately alters the relevant content in the “Thinking” part of the original positive example. Specifically, for every positive trajectory we generate five distinct negative counterparts, resulting in 10,000 positive and 50,000 negative trajectories for Softmax-DPO training.
@@ C.1 Model Initialization
As shown in Table 11, we initialize CogFlow with a 7B-scale pretrained MLLM backbone, Qwen2.5-VL-7B (Qwen2.5-VL). All parameters of the vision encoder, perception modules, and language decoder are updated during training without freezing any component.
@@ C.2 The Details of VPR
The Visual Parameterized Reward (VPR) metric quantifies perceptual accuracy within the parameter space of geometric primitives. As depicted in Figure 4, VPR first translates structured visual information into parametric representations. Critically, such structured inputs are constrained to three primitive types: points, lines, and circles, which constitute the foundational elements of geometric constructions alphageometry. For lines, parametric representations are derived by fitting straight-line models to the input data; for circles, parameters are extracted directly as center coordinates and radii. Subsequently, an optimal one-to-one correspondence $\mathcal{H}$ between predicted and ground-truth primitives is established via a multi-class Hungarian algorithm. This matching minimizes the total assignment cost $\mathcal{S}_{\mathrm{VPR}}$. The complete computational procedure is formalized in Algorithm 1.
@@ C.3 The Details of IntlzR Training
As shown in Table 11, the Knowledge Internalization Reward (IntlzR) model is trained on the MathCog-IntlzR subset containing 10k pairs (one positive and five negatives per pair). We initialize the reward model with Qwen2.5-VL-3B (Qwen2.5-VL) using the HuggingFace Transformers library, and train it for 3 epochs with a batch size of 64 and a learning rate of $7\times 10^{-6}$. It is worth noting that we adopt Softmax-DPO for training, with the objective function defined in Eq. 2, where the temperature parameter $\beta$ is fixed to 1.
@@ C.4 The Details of InfR
Accuracy Reward. The accuracy reward evaluates whether the response is correct. Following grpo, we assess correctness solely based on the answer field: the reward is 1 if the answer is correct and 0 otherwise.
@@ C.5 Details of SFT and RL Training
For RL, the training is conducted on the MathCog-RL subset (10k samples) for 1 epoch, with a learning rate of $1\times 10^{-6}$ and batch size of 16. For each input, we sample 8 candidate trajectories using temperature-controlled decoding. The rewards are computed by combining Synergistic Visual Rewards (SynVRs), Knowledge Internalization Reward (IntlzR), and Inference Reward (InfR), with their weights fixed to $(1,1,1)$. We implement reinforcement learning under the Verl (verl) framework on 16 NVIDIA A100 GPUs, with distributed acceleration provided by vLLM. Mixed-precision training is conducted in bfloat16, and gradient accumulation is employed to simulate larger effective batch sizes. We adopt DeepSpeed ZeRO-2 for memory-efficient optimization, and the KL penalty coefficient is set to $0.001$ to stabilize policy updates.
@@ C.6 Details of the Reasoning Drift Precision Metric
As illustrated in Figure 1, we assess the precision of reasoning drift: whether each reasoning step is faithfully grounded in the perceived structure. Concretely, a generated solution is decomposed into steps $\{s_{t}\}_{t=1}^{T}$. From each step, we extract the referenced visual cues (primitives and relations) $\mathcal{R}_{t}$ by GPT-5 gpt5, and from the perception parse $y^{\mathrm{w}*}$ we obtain the set of perceived elements $\mathcal{P}$. A step is deemed grounded if its visual claims are entailed by the perception, formalized by a consistency predicate, i.e., $D(s_{t})=1$. The precision of reasoning drift detection for a response is given by
@@ D.2 More Analysis on Synergistic Visual Rewards
Gemini-1.0-Pro (gemini) 47.7 48.3 35.1 50.5 65.8 42.5
GPT-4V (gpt4v) 49.9 43.1 50.5 57.5 65.2 38.0
GPT-4o (gpt4o) 63.8 - - - - -
Claude 3.5 Sonnet (claude-3-5) 67.7 - - - - -
Gemini-2.0-Flash (gemini) 73.4 - - - - -
Doubao-pro-1.5 (seed2025seed15thinkingadvancingsuperbreasoning) 79.5 77.7 88.9 86.0 82.3 62.0
LLaVA-1.5-7B (LLaVA) 20.0 22.7 7.7 11.8 26.6 33.0
LLaMA-Adapter-V2-7B (gao2023llamaadapterv2parameterefficientvisual) 23.9 21.2 25.5 11.3 32.3 31.8
TinyLLaVA-3B (zhou2024tinyllavaframeworksmallscalelarge) 26.7 20.4 19.7 23.1 44.9 31.8
LLaVA-1.5-13B (LLaVA) 26.8 19.7 20.2 18.8 45.6 36.9
mPLUG-Owl2-7B (ye2024mplug) 22.2 22.7 23.6 10.2 27.2 27.9
MiniGPT-v2-7B (chen2023minigptv2largelanguagemodel) 23.1 18.6 26.0 13.4 30.4 30.2
G-LLaVA-7B (gao2023gllavasolvinggeometricproblem) 25.1 19.1 48.7 3.6 25.0 28.7
VCAR (vcar) 33.7 30.9 34.6 38.7 37.3 28.5
SPHINX-Plus (sphinx) 36.7 54.6 16.4 23.1 41.8 43.0
SVE-Math-7B (SVE-Math-Qwen2.5-7B) 37.4 31.9 53.9 29.0 41.4 30.8
MultiMath-7B (MultiMath) 50.0 40.1 66.8 61.8 50.0 33.0
X-REASONER (liu2025xreasonergeneralizablereasoningmodalities) 69.0 - - - - -
VL-Rethinker-7B (VL-Rethinker) 73.7 - - - - -
ReVisual-R1 (chen2025advancingmultimodalreasoningoptimized) 73.1 - - - - -
WeThink (yang2025wethink) 70.9 - - - - -
Skywork-R1V-38B (Skyworkr1v2) 60.6 - - - - -
@@
DAPO DAPO w/o SynVRs, w/o IntlzR 56.1 45.8 4.1
@@
GRPO grpo w/o SynVRs, w/o IntlzR 56.7 47.6 4.2
@@ D.4 Analysis on Visual-Gated Policy Optimization
Table 13 presents an ablation analysis across different reinforcement learning algorithms (DAPO (DAPO), PPO (PPO), GRPO (grpo), and our VGPO) under consistent reward configurations based on the MathCog. Several clear trends emerge. First, the inclusion of visual perceptual reward (VPR) and Knowledge Internalization Reward (IntlzR) consistently improves performance across all methods, confirming their complementary roles in strengthening visual grounding. Second, GRPO outperforms both PPO and DAPO in all settings, highlighting its stability and effectiveness in geometry-oriented reasoning tasks. Most importantly, our proposed VGPO substantially surpasses all baselines, achieving 66.0%/56.2% accuracy on FlowVerse and MathVerse, respectively. This represents an improvement of over +6 points compared to GRPO with VPR and IntlzR, demonstrating the advantage of explicitly disentangling perception and reasoning optimization. Although VGPO incurs slightly higher training costs, the performance gain justifies the trade-off.
@@ D.6 Deeper insights into the mechanism of multi-stage reward structure
On the one hand, each reward is designed to target a distinct failure mode in visual mathematical reasoning: Firstly, the Synergistic Visual Rewards (SynVRs), which comprise VPR and VSR, are designed to address perceptual failures. Figure 5 presents an ablation of the SynVRs components. The results show that adding either VPR or VSR alone consistently improves performance, and the best results are obtained when both are enabled. Figure 15 further reports the perception F1 scores for different SynVRs variants, indicating that removing either reward degrades perception fidelity. Finally, the error-type analysis in Figure 9 demonstrates that SynVRs effectively reduce perception-related errors, thereby validating their effectiveness. Taken together, these results show that VPR and VSR provide complementary supervision signals: VPR enforces local geometric fidelity, while VSR encourages global perceptual coherence in terms of overall style and layout, and together they form trustworthy visual cues that serve as a robust foundation for effective visual mathematical reasoning, which is consistent with the conclusions reported in prior studies (DVLR; MathFlow). Secondly, the Knowledge Internalization Reward (IntlzR) is designed to alleviate reasoning drift by bridging the perception and reasoning stages, encouraging the model to produce structured, reasoning-ready outputs (i.e., knowledge-internalized representations (cognition4)) that provide a more reliable foundation for subsequent reasoning. The ablation study on IntlzR (Figure 6) shows that removing any single error type consistently degrades performance, indicating that each error type provides complementary supervision. The largest drops occur when excluding the omission/misbinding primitives or contradicting geometric constraints error types, highlighting that correctly binding primitives and respecting core geometric constraints are most critical for keeping reasoning tied to perception. In addition, the error-type analysis in Figure 9 further demonstrates that IntlzR effectively reduces Knowledge Internalization Errors, which in turn leads to improved overall performance. Thirdly, the Inference Reward (InfR) ensures task-level correctness and proper output structure at the final reasoning stage (grpo).
@@ D.10 Evaluating FG-CLIP for Similarity Reward
We further evaluate the FG-CLIP similarity module by replacing it with several alternative pre-trained models (e.g., MetaCLIP2 (metaclip)) of comparable capacity. Across these variants, the FG-CLIP (ViT-L-14) configuration consistently achieves the best overall performance on FlowVerse, indicating that FG-CLIP provides a stronger and more discriminative supervision signal for our VSR component than the alternative encoders

### v2

@@ Appendix B Additional Details of MathCog Dataset
Existing multimodal reasoning corpora provide abundant natural language annotations but rarely disentangle perception from reasoning (Zhang et al., 2025a; Gao et al., 2025). This gap makes it difficult to supervise models in a way that enforces accurate perception and strengthens their ability to internalize visual content. To address this, we construct the MathCog dataset, which explicitly separates the watching (perception) and thinking (reasoning) stages. Furthermore, to support the different training phases of CogFlow, we curate three tailored subsets: MathCog-SFT, MathCog-IntlzR, and MathCog-RL.
@@ B.2.1 Data Collection
To construct MathCog, we first collect a large pool of geometry-related visual math problems from existing corpora, including MAVIS (Zhang et al., 2025a), Geo170K (Gao et al., 2025), and LLaVA-CoT (Xu et al., 2025a). We then perform a careful filtering process to obtain a high-quality subset of 111,752 problems. Specifically, we remove (1) problems dominated by weak geometric relevance charts (e.g., histograms, pie charts) or artistic illustrations that deviate from geometric reasoning, and (2) diagrams with either excessively low resolution (<< 64) or overly high resolution (>> 2048), which would otherwise compromise model training consistency.
@@ B.2.2 MathCog-SFT and MathCog-RL Curation
To unify the representation, the extracted elements are normalized and remapped to a fixed coordinate system, following the normalization scheme in prior work (Wei et al., 2024):
@@ B.2.3 MathCog-IntlzR Curation
We first sample 10,000 examples as positive samples from MathCog-SFT for MathCog-IntlzR. As mentioned earlier, we categorize the errors into five types: (1) omitting or misbinding primitives, (2) introducing nonexistent facts, (3) contradicting geometric constraints, (4) invoking external theorems inappropriately, and (5) referring inconsistently to established elements. The negative data for MathCog-IntlzR is constructed based on these five error types. The curation of MathCog-IntlzR is illustrated at the bottom of the Figure 10. We first use an LLM, such as GPT-5 (OpenAI, 2025), to generate negative trajectories conditioned on each positive trajectory by explicitly instructing it to modify the corresponding “Thinking” part, as specified in the prompt shown in Table 7. The generated candidates are then manually checked. For example, for the Type 1 error: Omit or misbind primitives in Figure 11, we input the designed prompt into the LLM and obtain the variant “In any parallelogram, if the perimeter of triangle CDM is 5, …”, which deliberately alters the relevant content in the “Thinking” part of the original positive example. Specifically, for every positive trajectory we generate five distinct negative counterparts, resulting in 10,000 positive and 50,000 negative trajectories for Softmax-DPO training.
@@ C.1 Model Initialization
As shown in Table 11, we initialize CogFlow with a 7B-scale pretrained MLLM backbone, Qwen2.5-VL-7B (Bai et al., 2025). All parameters of the vision encoder, perception modules, and language decoder are updated during training without freezing any component.
@@ C.2 The Details of VPR
The Visual Parameterized Reward (VPR) metric quantifies perceptual accuracy within the parameter space of geometric primitives. As depicted in Figure 4, VPR first translates structured visual information into parametric representations. Critically, such structured inputs are constrained to three primitive types: points, lines, and circles, which constitute the foundational elements of geometric constructions Trinh et al. (2024). For lines, parametric representations are derived by fitting straight-line models to the input data; for circles, parameters are extracted directly as center coordinates and radii. Subsequently, an optimal one-to-one correspondence $\mathcal{H}$ between predicted and ground-truth primitives is established via a multi-class Hungarian algorithm. This matching minimizes the total assignment cost $\mathcal{S}_{\mathrm{VPR}}$. The complete computational procedure is formalized in Algorithm 1.
@@ C.3 The Details of IntlzR Training
As shown in Table 11, the Knowledge Internalization Reward (IntlzR) model is trained on the MathCog-IntlzR subset containing 10k pairs (one positive and five negatives per pair). We initialize the reward model with Qwen2.5-VL-3B (Bai et al., 2025) using the HuggingFace Transformers library, and train it for 3 epochs with a batch size of 64 and a learning rate of $7\times 10^{-6}$. It is worth noting that we adopt Softmax-DPO for training, with the objective function defined in Eq. 2, where the temperature parameter $\beta$ is fixed to 1.
@@ C.4 The Details of InfR
Accuracy Reward. The accuracy reward evaluates whether the response is correct. Following Shao et al. (2024), we assess correctness solely based on the answer field: the reward is 1 if the answer is correct and 0 otherwise.
@@ C.5 Details of SFT and RL Training
For RL, the training is conducted on the MathCog-RL subset (10k samples) for 1 epoch, with a learning rate of $1\times 10^{-6}$ and batch size of 16. For each input, we sample 8 candidate trajectories using temperature-controlled decoding. The rewards are computed by combining Synergistic Visual Rewards (SynVRs), Knowledge Internalization Reward (IntlzR), and Inference Reward (InfR), with their weights fixed to $(1,1,1)$. We implement reinforcement learning under the Verl (Sheng et al., 2025) framework on 16 NVIDIA A100 GPUs, with distributed acceleration provided by vLLM. Mixed-precision training is conducted in bfloat16, and gradient accumulation is employed to simulate larger effective batch sizes. We adopt DeepSpeed ZeRO-2 for memory-efficient optimization, and the KL penalty coefficient is set to $0.001$ to stabilize policy updates.
@@ C.6 Details of the Reasoning Drift Precision Metric
As illustrated in Figure 1, we assess the precision of reasoning drift: whether each reasoning step is faithfully grounded in the perceived structure. Concretely, a generated solution is decomposed into steps $\{s_{t}\}_{t=1}^{T}$. From each step, we extract the referenced visual cues (primitives and relations) $\mathcal{R}_{t}$ by GPT-5 OpenAI (2025), and from the perception parse $y^{\mathrm{w}*}$ we obtain the set of perceived elements $\mathcal{P}$. A step is deemed grounded if its visual claims are entailed by the perception, formalized by a consistency predicate, i.e., $D(s_{t})=1$. The precision of reasoning drift detection for a response is given by
@@ D.2 More Analysis on Synergistic Visual Rewards
Gemini-1.0-Pro (Team et al., 2025) 47.7 48.3 35.1 50.5 65.8 42.5
GPT-4V (OpenAI, 2023) 49.9 43.1 50.5 57.5 65.2 38.0
GPT-4o (OpenAI, 2024) 63.8 - - - - -
Claude 3.5 Sonnet (Anthropic, 2024) 67.7 - - - - -
Gemini-2.0-Flash (Team et al., 2025) 73.4 - - - - -
Doubao-pro-1.5 (Seed et al., 2025) 79.5 77.7 88.9 86.0 82.3 62.0
LLaVA-1.5-7B (Liu et al., 2024b) 20.0 22.7 7.7 11.8 26.6 33.0
LLaMA-Adapter-V2-7B (Gao et al., 2023) 23.9 21.2 25.5 11.3 32.3 31.8
TinyLLaVA-3B (Zhou et al., 2024) 26.7 20.4 19.7 23.1 44.9 31.8
LLaVA-1.5-13B (Liu et al., 2024b) 26.8 19.7 20.2 18.8 45.6 36.9
mPLUG-Owl2-7B (Ye et al., 2024) 22.2 22.7 23.6 10.2 27.2 27.9
MiniGPT-v2-7B (Chen et al., 2023a) 23.1 18.6 26.0 13.4 30.4 30.2
G-LLaVA-7B (Gao et al., 2025) 25.1 19.1 48.7 3.6 25.0 28.7
VCAR (Jia et al., 2024) 33.7 30.9 34.6 38.7 37.3 28.5
SPHINX-Plus (Liu et al., 2024a) 36.7 54.6 16.4 23.1 41.8 43.0
SVE-Math-7B (Zhang et al., 2025b) 37.4 31.9 53.9 29.0 41.4 30.8
MultiMath-7B (Peng et al., 2024a) 50.0 40.1 66.8 61.8 50.0 33.0
X-REASONER (Liu et al., 2025) 69.0 - - - - -
VL-Rethinker-7B (Wang et al., 2025a) 73.7 - - - - -
ReVisual-R1 (Chen et al., 2025c) 73.1 - - - - -
WeThink (Yang et al., 2025) 70.9 - - - - -
Skywork-R1V-38B (Wang et al., 2025c) 60.6 - - - - -
@@
DAPO Yu et al. (2025) w/o SynVRs, w/o IntlzR 56.1 45.8 4.1
@@
GRPO Shao et al. (2024) w/o SynVRs, w/o IntlzR 56.7 47.6 4.2
@@ D.4 Analysis on Visual-Gated Policy Optimization
Table 13 presents an ablation analysis across different reinforcement learning algorithms (DAPO (Yu et al., 2025), PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and our VGPO) under consistent reward configurations based on the MathCog. Several clear trends emerge. First, the inclusion of visual perceptual reward (VPR) and Knowledge Internalization Reward (IntlzR) consistently improves performance across all methods, confirming their complementary roles in strengthening visual grounding. Second, GRPO outperforms both PPO and DAPO in all settings, highlighting its stability and effectiveness in geometry-oriented reasoning tasks. Most importantly, our proposed VGPO substantially surpasses all baselines, achieving 66.0%/56.2% accuracy on FlowVerse and MathVerse, respectively. This represents an improvement of over +6 points compared to GRPO with VPR and IntlzR, demonstrating the advantage of explicitly disentangling perception and reasoning optimization. Although VGPO incurs slightly higher training costs, the performance gain justifies the trade-off.
@@ D.6 Deeper insights into the mechanism of multi-stage reward structure
On the one hand, each reward is designed to target a distinct failure mode in visual mathematical reasoning: Firstly, the Synergistic Visual Rewards (SynVRs), which comprise VPR and VSR, are designed to address perceptual failures. Figure 5 presents an ablation of the SynVRs components. The results show that adding either VPR or VSR alone consistently improves performance, and the best results are obtained when both are enabled. Figure 15 further reports the perception F1 scores for different SynVRs variants, indicating that removing either reward degrades perception fidelity. Finally, the error-type analysis in Figure 9 demonstrates that SynVRs effectively reduce perception-related errors, thereby validating their effectiveness. Taken together, these results show that VPR and VSR provide complementary supervision signals: VPR enforces local geometric fidelity, while VSR encourages global perceptual coherence in terms of overall style and layout, and together they form trustworthy visual cues that serve as a robust foundation for effective visual mathematical reasoning, which is consistent with the conclusions reported in prior studies (Guo et al., 2025c; Chen et al., 2025a). Secondly, the Knowledge Internalization Reward (IntlzR) is designed to alleviate reasoning drift by bridging the perception and reasoning stages, encouraging the model to produce structured, reasoning-ready outputs (i.e., knowledge-internalized representations (Ryan and Connell, 1989)) that provide a more reliable foundation for subsequent reasoning. The ablation study on IntlzR (Figure 6) shows that removing any single error type consistently degrades performance, indicating that each error type provides complementary supervision. The largest drops occur when excluding the omission/misbinding primitives or contradicting geometric constraints error types, highlighting that correctly binding primitives and respecting core geometric constraints are most critical for keeping reasoning tied to perception. In addition, the error-type analysis in Figure 9 further demonstrates that IntlzR effectively reduces Knowledge Internalization Errors, which in turn leads to improved overall performance. Thirdly, the Inference Reward (InfR) ensures task-level correctness and proper output structure at the final reasoning stage (Shao et al., 2024).
@@ D.10 Evaluating FG-CLIP for Similarity Reward
We further evaluate the FG-CLIP similarity module by replacing it with several alternative pre-trained models (e.g., MetaCLIP2 (Chuang et al., 2025)) of comparable capacity. Across these variants, the FG-CLIP (ViT-L-14) configuration consistently achieves the best overall performance on FlowVerse, indicating that FG-CLIP provides a stronger and more discriminative supervision signal for our VSR component than the alternative encoders